OpenStudy (anonymous):

how can you prove or disprove the scarecrow's statement of the pythagorean thereom?

OpenStudy (anonymous):

"The sum of the square roots of any two sides of an isosceles triangle is equal to the square root of the remaining side" always false

OpenStudy (anonymous):

But how do you know?

OpenStudy (anonymous):

i went to a talk at a math conference and the guy proved it i forget how, but i don't think it was that difficult

OpenStudy (anonymous):

in any case it is not pythagoras at all

OpenStudy (anonymous):

ok thankyhu!:)

OpenStudy (accessdenied):

If you say the segments creating the vertex have length a, and the base has length b. Just try writing the statement of each: sqrt(a) + sqrt(a) = sqrt(b) sqrt(a) + sqrt(b) = sqrt(a) It seems you end up with the base equal to zero and 2sqrt(a) = sqrt(b) = 0.